Find two numbers whose sum is 131 such that one number is 5 less than three times the other number
Answer by mananth(16946)

let the numbers be x & y
x+y =131....................1
...
x=3y-5
x-3y=5.......................2
..
multiply (1) by -1
we get
-x-y=-131
add this to (2)
-x+x-y-3y=-131+5
-4y=-126
/-4
y=-126/-4
y= 63/2
..
plug value of y in (1)
x+63/2= 131
x=131-63/2
x= 199/2
...
m.ananth@hotmail.ca